OpenStudy (anonymous):

Solve the inequality \[(\frac{4}{3x+2}) ge (0)\] and express the solution as an interval.

OpenStudy (anonymous):

\[\frac{4}{3x+2} \ge 0\]

OpenStudy (anonymous):

changes sign where the denominator is 0, namely at \(-\frac{2}{3}\)

OpenStudy (anonymous):

positive if \(x>-\frac{2}{3}\) negative otherwise

OpenStudy (anonymous):

So \[x \ge -\frac{2}{3}\]

OpenStudy (anonymous):

I mean yea not equal to...

OpenStudy (anonymous):

make sure to answer with a strict inequality \(>\) rather than \(\geq\) because the denominator cannot be zero

OpenStudy (anonymous):

Okay, thank you